OpenStudy (anonymous):

ok so here is a new on Which line has a slope of 2 and goes through (2, 10)?

OpenStudy (anonymous):

the one that looks like \[y-10=2(x-2)\] when you solve for \(y\)

OpenStudy (anonymous):

an application of the points slope formula \[y-y_1=m(x-x_1)\]